OpenStudy (volleyballlover55):

Find an irrational number that is between 5.2 and 5.5. Explain why it is irrational. Include the decimal approximation of the irrational number to the nearest hundredth.

OpenStudy (volleyballlover55):

@jim_thompson5910

OpenStudy (volleyballlover55):

yay! your back :)

jimthompson5910 (jim_thompson5910):

let x be the unknown number x is between 5.2 and 5.5, so, \[\Large 5.2 < x < 5.5\] \[\Large (5.2)^2 < x^2 < (5.5)^2\] \[\Large 27.04 < x^2 < 30.25\] agreed so far? or no?

OpenStudy (volleyballlover55):

yes

jimthompson5910 (jim_thompson5910):

what whole number is between 27.04 and 30.25?

OpenStudy (volleyballlover55):

28

jimthompson5910 (jim_thompson5910):

what is the square root of 28?

OpenStudy (volleyballlover55):

5.29150262213

jimthompson5910 (jim_thompson5910):

do you agree that 5.29150262213 is between 5.2 and 5.5 ?

OpenStudy (volleyballlover55):

yes

jimthompson5910 (jim_thompson5910):

is 5.29150262213 rational ? or irrational?

OpenStudy (volleyballlover55):

irrational

jimthompson5910 (jim_thompson5910):

yep because there is no pattern that repeats. It seems like it's random so 5.29150262213 can't be written as a fraction of two whole numbers

OpenStudy (volleyballlover55):

right!

jimthompson5910 (jim_thompson5910):

so by extension, \(\Large \sqrt{28} \) is also irrational

OpenStudy (volleyballlover55):

ok so how would we put that answer in a sentence?

jimthompson5910 (jim_thompson5910):

just write what I posted but in your own words

OpenStudy (volleyballlover55):

um is this a good answer? An irrational number between 5.2 and 5.5 is 5.29150262213. It is irrational because there is no pattern that repeats and it cant be written as a fraction of two whole numbers.

jimthompson5910 (jim_thompson5910):

yes that works

OpenStudy (volleyballlover55):

awesome! next! :)

OpenStudy (volleyballlover55):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

tell me what you're thinking

OpenStudy (volleyballlover55):

i think conner is correct

jimthompson5910 (jim_thompson5910):

why

OpenStudy (volleyballlover55):

i dont know.. lol

OpenStudy (volleyballlover55):

can you tell me..

jimthompson5910 (jim_thompson5910):

this might help http://www.onlinemathlearning.com/image-files/387x248xari121.gif.pagespeed.ic.vFV2Lan_oa.png

OpenStudy (volleyballlover55):

ok do you know the answer?

jimthompson5910 (jim_thompson5910):

yes but you should look at the link

OpenStudy (volleyballlover55):

ok i saw it :)

jimthompson5910 (jim_thompson5910):

what is 2^7 times 2^9 equal to write it in exponential form

OpenStudy (volleyballlover55):

65536

jimthompson5910 (jim_thompson5910):

write it in exponential form

jimthompson5910 (jim_thompson5910):

example: 32 in exponential form is 2^5

OpenStudy (volleyballlover55):

i what 65536 is though in exponential form

OpenStudy (volleyballlover55):

i dont know*

jimthompson5910 (jim_thompson5910):

hint: rule 1 in that link

OpenStudy (volleyballlover55):

i dont get it

jimthompson5910 (jim_thompson5910):

example: 7^5 times 7^8 = 7^(5+8) = 7^13

OpenStudy (volleyballlover55):

but 65536 is a long number.. can you wirte it in exponential form for me..

OpenStudy (volleyballlover55):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

use my last example to do 2^7 times 2^9